### 12 Dec 2023 v1

@@ Abstract
Strong Branching (SB) is a cornerstone of all modern branching rules used in the Branch-and-Bound (BNB) algorithm, which is at the center of Mixed-Integer Programming solvers. In its full form, SB evaluates all variables to branch on and then selects the one producing the best relaxation, leading to small trees, but high runtimes. State-of-the-art branching rules therefore use SB with working limits to achieve both small enough trees and short run times. So far, these working limits have been established empirically.
In this paper, we introduce a theoretical approach to guide how much SB to use at each node within the BNB. We start by representing the BNB algorithm as an abstract stochastic tree model where the geometric mean of the dual gains of all variables follows a given probability distribution. This model allows us to relate expected dual gains to tree sizes and explicitly compare the cost of sampling an additional SB candidate with the reward in expected tree size reduction. We then leverage the insight from the abstract model to design a new stopping criterion for SB, which fits a distribution to the dual gains and, at each node, dynamically continues or interrupts SB. This algorithm, which we refer to as Probabilistic Lookahead Strong Branching, improves both the tree size and runtime over MIPLIB instances, providing evidence that the method not only changes the amount of SB, but allocates it better.
@@ 1 Introduction
The core of most MIP solvers is the Branch-and-Bound algorithm (BNB) [1]. It recursively partitions the solution space of the MIP into subproblems and computes their continuous Linear Programming (LP) relaxations. If the LP solution to a subproblem is integral, i.e., it is a solution to the MIP, then the best MIP solution is updated. If the optimal value of the LP relaxation of a subproblem is not better than the current best integer solution known for the MIP, this subproblem can be pruned. Otherwise, the subproblem is further divided. We refer interested readers to [11] for a recent overview on MIP and to [1] for the algorithmic aspects of MIP solving.
@@
The current default branching algorithm in SCIP is an extension of reliability [3] and hybrid branching [2]. It uses in particular strong branching calls for variables that are considered unreliable (i.e., which have not been branched on a certain number of times) and then relies on a weighted score dominated by pseudocosts for reliable variables. Even when all candidates are still unreliable – e.g., at the root node – evaluating all of them would be prohibitively expensive, solving twice as many LPs as there are candidates. If the branching rule calls strong branching to evaluate candidates, a lookahead strategy is used: if a certain number of successive variables are evaluated without any improvement over the best candidate, the procedure stops and returns the current best. This maximum lookahead value is essentially proportional to a fixed parameter $L$. Conceptually, this problem can be viewed as a generalized secretary problem with interview cost [15, 6] or a variant of Pandora’s box problem [9, 24], in which indistinct candidates with hidden rewards have to be evaluated sequentially to maximize a net difference between the selected candidate’s reward and the total interview costs. We will refer to the number of unsuccessful successive candidates as the lookahead value throughout this paper and will propose a novel approach to solve the selection and stopping problem equipped with a dynamic lookahead replacing the static $L$ value.
A probabilistic model of variable dual gains was proposed in [17] in order to determine whether a given variable is unreliable and to choose which one to select for branching. More precisely, the dual gains of any variable are modeled as random variables following a normal distribution (i.i.d. across all nodes), and a parametric statistical test is used to determine a probability that a variable can yield a greater gain than the incumbent. In contrast, our probabilistic model represents the dual gains of all variables at the current node as drawn from a distribution, making our approach applicable even when some variables have not been sampled yet, and therefore from the very first nodes of the tree.
Contributions. We define an abstract model of the Branch-and-Bound algorithm, based on the Multi-Variable Branching (MVB) model of [19], incorporating unknown variable gains akin to Pandora’s box problem and a particular form of tree restart. We coin the new model as Pandora’s MVB, or PVB. We determine additional assumptions under which strong branching should continue or stop and branch on the best candidate found so far, defining a stopping criterion for our Probabilistic Lookahead Strong Branching (PL-SB) algorithm.
@@ 3 Pandora’s MVB: A Stochastic Abstract Model for Branch-and-Bound
Therefore, we start by constructing an abstract BNB model extending the approach of [19], which has proven to capture important aspects of BNB well enough to improve performance in SCIP. However, all the models proposed in [19] suppose that the dual gains are fixed throughout the tree and known in advance, which renders an algorithm like SB completely superfluous in this setting.
For this reason, we design in this section an extended abstract model in which the dual gains are fixed, but not known until the corresponding variable is branched on. Based on this model, we design in Section 4 a stopping criterion for SB based on the likelihood that additional strong branching iterations provide a better branching candidate, as measured by the tree size.
@@ 3.1 Background on BNB Abstract Modeling
where $x_{j}$ denotes the variable branched on at node $h$.
One of the problems defined using this abstract model is the Multiple Variable Branching (MVB) problem. The MVB model is a simplified version of the branch and bound tree, where $n$ variables are given in input, and where one variable $x_{j}$ with gains $(l_{j},r_{j})$ must be branched on at each node to build a tree that closes a gap $G$. In MVB, one variable can be branched on multiple times on a path from the root to a leaf. However, because of the assumptions of MVB, some aspects of a modern implementation of a BNB, like SB, cannot be modeled. Indeed, in a BNB in which the dual gains of the variables would be fixed and known in advance as they are in MVB, candidate sampling as performed by SB would not be relevant. Therefore, in Section 3.2, we introduce an abstract model in which the dual gains of the variables are unknown.
@@ 3.2 An Abstract BNB Model with Stochastic Variable Gains
We will represent SB in an abstract BNB model using restarts, a common technique from constraint and mixed-integer programming. We define a restart as the removal of all nodes from the tree, with two important properties: (a) the nodes evaluated before a restart are still accounted for in the total cost of the tree, and (b) the observed dual gains remain valid after a restart. In this setting, a SB iteration is a restart after branching on one variable, so that its dual gain can be determined without actually branching on it in the final tree that closes the gap $G$. We can now define the PVB problem on our abstract BNB model:
@@
Given $n\in\mathbb{Z}_{+}$ variables with random dual gains in $\mathbb{Z}_{+}^{2}$ that follow their own distribution, a target gap $G\in\mathbb{Z}_{+}$, and $k\in\mathbb{Z}_{+}$, can a BNB tree that closes the gap $G$ be built using at most $k$ nodes, using each variable as many times as needed, where the random dual gains of a variable are revealed when it is first branched on, and where restarts are allowed?
If we did not incorporate restarts in our model, the children of the root node would be created by branching on a variable without any information on its dual gains or that of other variables, which is very unlike what current state-of-the-art branching rules do. Note that one can effectively simulate SB in PVB by restarting after branching on one variable, and finally build the final tree without restarting.
@@ 4 Probabilistic Lookahead: a Stopping Criterion for Strong Branching on Pandora’s MVB
Section 2 shows the potential of more selectively using SB in MIP solvers. We now design a SB strategy for PVB, and compare it to SCIP’s default rule on PVB instances. Simulation experiments on instances of the new abstract model show that the new criterion is substantially better in total LP evaluations than the fixed maximum lookahead stopping criterion used in SCIP.
@@ 4.1 Necessary Assumptions
The smallest tree that closes a gap $G$ by repeatedly branching on the same variable is referred to as SVB tree, in line with [19]. Of course, no variable is known at the beginning, therefore the goal is to decide when the best variable found so far is “good enough” compared to the variables that still have unknown dual gains. In order to do so, we first make a number of explicit assumptions:
@@
    The size of an SVB tree that closes $G$ by only branching on variable $(g_{j},g_{j})$ is equal to the size of the tree that only branches on $(l_{j},r_{j})$, where $g_{j}=\sqrt{l_{j}r_{j}}$ is the geometric mean of variable $x_{j}$.
@@
We focus on the geometric mean of a variable because its square, the product, is a state-of-the-art scoring function for variables [1]. Hence, it has shown that it is a good practical heuristic for ranking branching candidates. We now use the assumptions above to design a stopping criterion for SB.
@@ 4.3 Evaluating the Probabilistic Lookahead on Pandora’s MVB
Because we observed that the performance of these stopping criteria is similar relative to the others on all instances, here we only present results on the Trento1 MIPLIB instance. For each gap value $G$, we conduct 1000 experiments. The average overall number of nodes and number of nodes used for SB are presented in Table 2. We also conducted random branching experiments, computing the expected tree size over all fractional variables. However, the average random tree size turned out to be essentially infinite, because some variables have quasi-zero dual gains.
@@ 5 The Geometric Mean of Dual Gains Follows a Mixed Distribution
Before presenting the algorithm we design for strong branching, we assess the accuracy of Assumption A2, namely, that there is a probability distribution that captures well the geometric mean of dual gains. In practice, the dual bound change when branching on a given variable depends heavily on the particular structure of the problem. Fixing decision variables that impact the rest of the problem structure – e.g., design variables in a network design problem – will result in structural changes that are likely to change the relaxation bound a lot more than variables that have a more “local” influence on – e.g., routing variables in the same network design example. We compare the distribution of mean dual gains for the MIPLIB 2017 benchmark instances at the root node with different distributions fitted with their maximum likelihood estimators. Three requirements for the distribution we fit are having a zero probability on negative gains, allowing for a mass point at zero, and having a support that is a proper superset of the interval $\left[0,g_{\max}\right]$, since we want to estimate the probability of a new sample having a value greater than the current maximum mean dual gain $g_{\max}$, which would be zero if the support is equal to the interval, e.g., with the uniform or empirical cumulative distributions. The reason for the mass point at zero is the dual gain often being zero for many variables, i.e., the LP optimum after branching remains on the original optimal face. To accommodate this characteristic, we model dual gains with a mixed continuous-discrete distribution:
@@ 8.0.1 Acknowledgements
Research reported in this paper was partially supported through the Research Campus Modal funded by the German Federal Ministry of Education and Research (fund numbers 05M14ZAM, 05M20ZBM).

### 23 Jan 2024 v2

@@ Abstract
Strong Branching (SB) is a cornerstone of all modern branching rules used in the Branch-and-Bound (BnB) algorithm, which is at the center of Mixed-Integer Programming solvers. In its full form, SB evaluates all variables to branch on and then selects the one producing the best relaxation, leading to small trees, but high runtimes. State-of-the-art branching rules therefore use SB with working limits to achieve both small enough trees and short run times. So far, these working limits have been established empirically. In this paper, we introduce a theoretical approach to guide how much SB to use at each node within the BnB. We first define an abstract stochastic tree model of the BnB algorithm where the geometric mean dual gains of all variables follow a given probability distribution. This model allows us to relate expected dual gains to tree sizes and explicitly compare the cost of sampling an additional SB candidate with the reward in expected tree size reduction. We then leverage the insight from the abstract model to design a new stopping criterion for SB, which fits a distribution to the dual gains and, at each node, dynamically continues or interrupts SB. This algorithm, which we refer to as Probabilistic Lookahead Strong Branching, improves both the tree size and runtime over MIPLIB instances, providing evidence that the method not only changes the amount of SB, but allocates it better.
@@ 1 Introduction
The core of most MIP solvers is the Branch-and-Bound algorithm (BnB) [1]. It recursively partitions the solution space of the MIP into subproblems and computes their continuous Linear Programming (LP) relaxations. If the LP solution to a subproblem is integral, i.e., it is a solution to the MIP, then the best MIP solution is updated. If the optimal value of the LP relaxation of a subproblem is not better than the current best integer solution known for the MIP, this subproblem can be pruned. Otherwise, the subproblem is further divided. We refer interested readers to Conforti et al. [11] for a recent overview on MIP and to Achterberg [1] for the algorithmic aspects of MIP solving.
@@
The current default branching algorithm in SCIP is an extension of reliability [3] and hybrid branching [2]. It uses in particular strong branching calls for variables that are considered unreliable (i.e., which have not been branched on a certain number of times) and then relies on a weighted score dominated by pseudocosts for reliable variables. Even when all candidates are still unreliable – e.g., at the root node – evaluating all of them would be prohibitively expensive, solving twice as many LPs as there are candidates. If the branching rule calls strong branching to evaluate candidates, a lookahead strategy is used: if a certain number of successive variables are evaluated without any improvement over the best candidate, the procedure stops and returns the current best. This maximum lookahead value is essentially proportional to a fixed parameter $L$. Conceptually, this problem can be viewed as a generalized secretary problem with interview cost [15, 6] or a variant of Pandora’s box problem [9, 24], in which indistinct candidates with hidden rewards have to be evaluated sequentially to maximize a net difference between the selected candidate’s reward and the total interview costs. We will refer to the number of unsuccessful successive candidates as the lookahead value throughout this paper and will propose a novel stopping criterion equipped with a dynamic lookahead replacing the static $L$ value.
A probabilistic model of variable dual gains was proposed in Hendel [17] in order to determine whether a given variable is unreliable and to choose which one to select for branching. More precisely, the dual gains of any variable are modeled as random variables following a normal distribution (i.i.d. across all nodes), and a parametric statistical test is used to determine a probability that a variable can yield a greater gain than the incumbent. This per-variable distribution can only provide guidance deep in the tree when a variable has been branched on several times to offer statistically significant inference. In contrast, our probabilistic model represents the dual gains of all variables at the current node as drawn from a distribution, making our approach applicable even when some variables have not been sampled yet, and therefore from the very first nodes of the tree. Furthermore, our model can be adapted to any distribution for which the cumulative function can be evaluated, alleviating the need for restricting assumptions such as normality.
Contributions. We define an abstract model of the Branch-and-Bound algorithm, based on the Multi-Variable Branching (MVB) model of Le Bodic and Nemhauser [19], incorporating unknown variable gains akin to Pandora’s box problem and a particular form of tree restart. We coin the new model as Pandora’s MVB, or PVB. We determine additional assumptions under which strong branching should continue or stop and branch on the best candidate found so far, defining a stopping criterion for our Probabilistic Lookahead Strong Branching (PL-SB) algorithm.
@@ 3 Pandora’s MVB: A Stochastic Abstract Model for Branch-and-Bound
Therefore, we start by constructing an abstract BnB model extending the approach of Le Bodic and Nemhauser [19], which has proven to capture important aspects of BnB well enough to improve performance in SCIP. However, all the models proposed in Le Bodic and Nemhauser [19] suppose that the dual gains are fixed throughout the tree and known in advance, which renders an algorithm like SB completely superfluous in this setting.
For this reason, we design in this section an extended abstract model in which the dual gains are fixed, but not known until the corresponding variable is branched on. This is not true in general. However, MIP solvers internally estimated the dual gains of variables by a pair of values, the pseudocosts. This assumption is useful to analyze the resulting abstract models. Based on this model, we design in Section 4 a stopping criterion for SB based on the likelihood that additional strong branching iterations provide a better branching candidate, as measured by the tree size.
@@ 3.1 Background on BnB Abstract Modeling
where node $h$ is the parent of $i$, if there is one, $G_{h}$ is the gap at $h$, and $(l_{j},r_{j})$ is the pair of gains of the variable branched on at $h$.
One of the problems defined using this abstract model is the Multiple Variable Branching (MVB) problem. The MVB model is a simplified version of the branch and bound tree, where $n$ variables are given in input, and where one variable $x_{j}$ with gains $(l_{j},r_{j})$ must be branched on at each node to build a tree that closes a gap $G$. In MVB, one variable can be branched on multiple times on a path from the root to a leaf. However, because of the simplicity of MVB, some aspects of a modern implementation of a BnB, such as SB, cannot be modeled. Indeed, in a BnB in which the dual gains of the variables would be fixed and known in advance as they are in MVB, information discovery from candidate sampling as performed by SB would not be relevant. Therefore, in Section 3.2, we introduce an abstract model in which the dual gains of the variables are unknown.
@@ 3.2 An Abstract BnB Model with Stochastic Variable Gains
We will represent SB in an abstract BnB model using restarts, a common technique from constraint and mixed-integer programming. We define a restart as the removal of all nodes from the tree, with two important properties: (a) the nodes evaluated before a restart are still accounted for in the total cost of the tree, and (b) the observed dual gains remain valid after a restart. Performing SB therefore consists in branching on one variable, recording its dual gain, restarting the tree, and iterating. We can now define the PVB problem on our abstract BnB model:
@@
Given $n\in\mathbb{Z}_{+}$ variables with random dual gains in $\mathbb{R}_{+}^{2}$ that follow their own distribution, a target gap $G\in\mathbb{R}_{+}$, and $k\in\mathbb{Z}_{+}$, can a BnB tree that closes the gap $G$ be built using at most $k$ nodes, using each variable as many times as needed, where the random dual gains of a variable are revealed when it is first branched on, and where restarts are allowed?
Note that one can effectively simulate SB in PVB by restarting after branching on one variable, and finally build the final tree without restarting.
@@ 4 Probabilistic Lookahead: a Stopping Criterion for Strong Branching on Pandora’s MVB
Section 2 shows the potential of more selectively using SB in MIP solvers. In this section, we design a SB strategy for PVB, and we compare it to SCIP’s default rule on PVB instances. Simulation experiments on instances of the new abstract model show that the new criterion is substantially better in total LP evaluations than the fixed maximum lookahead stopping criterion used in SCIP.
@@ 4.1 Necessary Assumptions
The smallest tree that closes a gap $G$ by repeatedly branching on a single variable is referred to as SVB tree, in line with [19]. In order to establish a stopping criterion that is grounded in theory, we first make a number of explicit assumptions:
@@
    The size of an SVB tree that closes $G$ by only branching on variable with gains $(g_{j},g_{j})$ is equal to the size of the tree that only branches on a variable with gains $(l_{j},r_{j})$, where $g_{j}=\sqrt{l_{j}r_{j}}$ is the geometric mean of dual gains of variable $x_{j}$.
@@
We focus on the geometric mean of a variable because its square, the product, is a state-of-the-art scoring function for variables [1]. Hence, it has been shown that it is a good practical heuristic for ranking branching candidates. We now use the assumptions above to design a stopping criterion for SB.
@@ 4.3 Evaluating the Probabilistic Lookahead on Pandora’s MVB
Because we observed that the performance of these stopping criteria is similar relative to the others on all instances, here we only present results on the Trento1 MIPLIB instance. For each gap value $G$, we conduct 1000 experiments. The average overall number of nodes and number of nodes used for SB are presented in Table 2.
@@ 5 The Geometric Mean of Dual Gains Follows a Mixed Distribution
Before presenting the algorithm we design for strong branching, we assess the accuracy of Assumption A2, namely, that there is a probability distribution that captures well the geometric mean of dual gains. In practice, the dual bound changes when branching on a given variable depend heavily on the particular structure of the problem. Fixing decision variables that impact the rest of the problem structure – e.g., design variables in a network design problem – will result in structural changes that are likely to change the relaxation bound a lot more than variables that have a more “local” influence on – e.g., routing variables in the same network design example. We compare the distribution of mean dual gains for the MIPLIB 2017 benchmark instances at the root node with different distributions fitted with their maximum likelihood estimators. Three requirements for the distribution we fit are having a zero probability on negative gains, allowing for a mass point at zero, and having a support that is a proper superset of the interval $\left[0,g_{\max}\right]$, since we want to estimate the probability of a new sample having a value greater than the current maximum mean dual gain $g_{\max}$, which would be zero if the support is equal to the interval, e.g., with the uniform or empirical cumulative distributions. The reason for the mass point at zero is the dual gain often being zero for many variables, i.e., the LP optimum after branching remains on the original optimal face. To accommodate this characteristic, we model dual gains with a mixed continuous-discrete distribution:
@@ 8.0.1 Acknowledgements
Research reported in this paper was partially supported through the Research Campus Modal funded by the German Federal Ministry of Education and Research (fund numbers 05M14ZAM, 05M20ZBM). We thank Nicolas Gast and Bruno Gaujal for discussions on online decision-making including the Pandora’s box problem.